TELL WEATHER THE RELATIONSHIP BETWEEN X AND Y IS LINEAR,EXPONENTIAL,OR NEITHER AND EXPLAIN THE ANSWER- X-0,1,2,3,4,5 -Y-1,1/4,1/16,1/64,1/256,1/1024 ?
Answer by stanbon(75887)

Y=(1/4)^X
Y(0)=(1/4)^0=1
Y(3)=(1/4)^3=1/64
etc.
This is an exponential equation.
Cheers,
Stan H.